# Path-based packing of icosahedral shells into multi-component aggregates

### v1

@@ Methods
Generalised anti-Mackay (AM) shells are non-closed-packed, achiral icosahedral shells. Each AM shell is identified univocally by the two AM indexes $p$ and $q$: $p$ is the number of particles on the side of the closed-packed equilateral triangle concentric to the icosahedral facet; $q$ is the number of particles on each median of the icosahedral facet, between the vertex of the icosahedral facet and the vertex of the inner triangle (see Fig. 3A in the main text). As for the CK shells, simple geometric considerations allow us to identify the rules for packing AM shells into icosahedra, and to associate the optimal size mismatch to each possible interface between nearby shells of different kind.
Here we start by calculating the number of particle in a AM shell, as a function of the two AM indexes $p$ and $q$. It is useful to divide the particles of the icosahedral facet into different groups, as shown in Fig. SS2:
@@
so that each AM shell can be put in relation with a CK shell of the same size. More precisely, the AM shell can transform into the corresponding CK shell by some concerted rotations of the particles, with different rotation axes (perpendicular to the planes of the icosahedral facets) and different rotation angles and, viceversa, CK shells can transform into AM ones. In Figs. 4, B and C of the main text we have shown two examples:
@@
The edge of the AM shell with unit inter-particle distance and AM indexes $p$, $q$ can be easily calculated by looking at Fig. SS2. It corresponds to the side of the inner triangle, which is $p-1$, plus twice the projection of the segment between the vertices of the facet and of the inner triangle (which is $q+1$ long) on the facet edge:
@@
Values are reported in Fig. 4D of the main text.
@@
The data on the optimal mismatch in Lennard-Jones and Morse clusters have been obtained by calculating the energy, after local minimization, as a function of the mismatch for the different types of structures, and looking for the mismatch at which the energy is minimum. Two examples are reported in Fig. SS3, where we consider MC, AM1 and Ch1 shells on MC cores. The MC structures are better stabilised in a low-mismatch range. The AM1 shells find their optimal mismatch in an intermediate range and after a threshold mismatch they become unstable, i.e. they are not even local minima. On the contrary, the Ch1 structures are stable in a high-mismatch range and become unstable below a threshold mismatch. In the cases reported in Fig. SS3 the range of mismatch values in which both AM1 and Ch1 structures are local minima is very narrow.
@@
Values of the optimal mismatch are reported in Fig. SS4. As in the case of chiral shells (see Fig. 2D in the main text), the optimal mismatch is larger for the class-changing step $(1,k)^{*}\to(2,k)^{*}$. This may lead to the conclusion that the outer shell should be preferentially of AM1 type (i.e. $(1,k+1)^{*}$) if the mismatch is zero or low. But this is not the case, as we show in Fig. SS5. In fact the mismatch range in which the configuration with two AM1 shells is a local minimum is very narrow, and even when locally stable, its energy is quite high. Outside that range, the outer shell spontaneously reconstructs upon local minimization, as shown in Fig. SS5, A and B. This is due to the very unfavourable coordination of particles on the facet edges of the outer shell, which sit on only two nearest neighbours of the shell below (this point will be further discussed in the last Section). Both reconstructions improve that coordination by displacing atoms on fourfold sites of the substrate. In Fig. SS5C we calculate the energies of unreconstructed and reconstructed configurations of Lennard-Jones clusters. The reconstruction of Fig. SS5B is more favourable, as it displaces a larger number of particles on sites of higher coordination with the substrate. Due to the instability of the configuration with two AM1 shells, the natural growth sequences lead to the formation of an outer AM2 shell, as shown in Fig. 4E of the main text and discussed in Section S3.
Finally, we note that the configuration with two Ch1 shells, which has the same number of particles, is much more favourable in a wide range of mismatch, as shown in Fig. SS5C. In fact, from the results in Fig. SS5C it turns out also that the stability range of two Ch1 shells is enlarged with respect to that of one Ch1 shell (see Fig. SS3A).
@@
In Tables S1-3 we report the data for the atomic pair exchanges shown in Fig. SS6 for different types of alkali metal clusters. DFT [48] and Gupta potential [49] results are reported. The results have been obtained as explained in the Methods section of the main text. Here we recall that the energy differences between the configuration after and before the exchange are related to locally minimized configurations.
All structures contain a core made of (0,0) and (0,1) MC shells, upon which further shells are added, either of Ch1 class (the (1,1) BG shell and the (1,2) chiral shell) or of Ch0 class (the (0,2) and (0,3) MC shells). In the latter case, standard Mackay icosahedra are obtained. According to the estimates of the optimal mismatch values, for Na@K, Na@Rb and Na@K@Rb we expect that the clusters with Ch1 shells are more stable than the Mackay clusters. This is confirmed by the data reported in Tables S1-3, where it is shown that Mackay clusters are energetically unstable upon atomic pair exchanges, while the clusters with Ch1 shells are stable. We note that DFT and Gupta results are in good agreement, which supports the validity of the Gupta model for these clusters.
Appendix S3 Growth simulations
Here we present additional results on the simulations of the growth after deposition of atoms one-by-one on a preformed seed.[39, 22] These simulations show a few more examples of growth sequences that can be represented by paths in the hexagonal lattice.
In Fig. SS7, three growth sequences for binary and ternary alkali metal clusters are shown. In Fig. SS7A, we show the result of the deposition of Cs atoms on a Na@K seed. The seed is terminated by a (1,3) Ch1 shell of K atoms. The deposition of Cs atoms, whose mismatch with K is of 0.154, lead to the formation of a Ch2 (2,3) shell, i.e. to a change of class caused by size mismatch. Cs and K atoms show some tendency towards intermixing,[50] so that there are a few exchanges of K and Cs atoms during growth that bring some K atoms to the cluster surface. In Fig. SS7B, we report the deposition of K atoms on a Na@K cluster terminated by a K (1,1) BG shell. This is a case of deposition without mismatch, therefore the growth continues within the Ch1 class as expected. In Fig. SS7C a more complex growth sequence is shown. The seed is Na@K terminated by a Ch1 (1,2) shell of K atoms. Rb atoms are deposited on top of this seed. Since the mismatch between Rb and K is rather small (0.072) one would expect that growth continues within the Ch1 class. However this is not the case, because the (2,2) BG shell of Ch2 class is formed, followed then by a (2,3) Ch2 shell. However this is not the end of the story, because upon further Rb atom deposition there is first a sudden transformation of the subsurface shell from (2,2) to (1,3) and then of the (2,3) shell from (2,3) to (1,4). These transformation take place by sudden collective reshaping processes, of the type frequently observed in metal clusters.[22] The final result is as if the cluster would have grown into the Ch1 class from the beginning, thus restoring the natural growth path as expected on the basis of the small mismatch between K and Rb atoms.
In Fig. SS8 data on the growth of Ag atoms on Ni and Co Mackay seeds of 147 and 309 atoms are shown. The large mismatch (0.161 for Ag on Ni and 0.156 for Ag on Co) is close to optimal for the growth of Ch1 shells on seeds in this size range; this kind of growth is indeed observed in the growth simulations.
In Fig. SS9 the growth of Ag atoms on a Mackay seeds of size 309 is shown. This cluster is terminated by a (0,4) shell. For the growth of the first Ag shell, the mismatch between Ag and Cu (0.129) is closer to the optimal one of an AM1 shell. The AM1 (1,4) shell is indeed grown, with very few defects. Then the growth on the AM1 shell continues by a class-changing path by forming defective AM shells up to the AM3 (1,4) shell. Upon further deposition, there is a sudden transformation of all Ag shells to the Ch1 arrangement, up to the (1,6); this configuration is energetically more favourable.
In Fig. SS10A, the growth of Pd atoms on a Mackay Ni core of size 147 is shown. The core si terminated by a (0,3) shell. The mismatch between Pd and Ni is 0.104, so that an AM1 (1,3) shell is expected to grow; which is indeed the case. Upon depositing Ag on the cluster terminated by the (1,3) Pd shell, a (2,3) AM2 shell is formed, due to the moderate mismatch between Ag and Pd, which is 0.051. The cluster further grows by forming a (3,3) BG shell, which is then covered by a (3,4) Ch3 shell (see Fig. SS10B).
In Fig. SS11 the growth of Ag on imperfect Mackay Ni seeds is shown. These seeds present an incomplete Mackay external shell. In spite of the defective character of the seed, Ch1 shells are formed. These shell are slightly distorted because of the asymmetries in the starting core, but they very much resemble the perfect Ch1 shells. This shows that the growth mode of chiral shell is quite robust against imperfections.
In Fig. SS12 the growth mechanism of AM shells one on top of the other is discussed. This mechanism is dominated by the trapping of atoms in fourfold adsorption sites on the surface, which naturally leads to class changes at every step in the hexagonal plane.
Figure S1: Implementation of the CK construction.
Figure S2: Facet of an anti-Mackay shell.
Figure S3: Stability of MC, AM1 and Ch1 shells in Lennard-Jones clusters.
Figure S4: Optimal mismatch for two AM shells on a MC core.
Figure S5: Stability of two AM shells on a MC core.
Figure S6: Clusters of the DFT calculations.
@@
The structures considered in the calculations are of two types, with either $i=3$ or $i=4$ shells. In this figure we show only those containing BH and Ch1 shells, while the analogous structures containing 3 or 4 MC shells are not shown. The first structure ($i=3$) contains the central atom (0,0), the (0,1) MC shell, and the (1,1) BG shell, corresponding to the left and middle panels in A, for a total of 45 atoms. The second structure ($i=4$) contains in addition the (1,2) Ch1 shell, as shown in the right panel of A, for a total of 117 atoms. In B-E the different types of atom pair swaps are shown: in B and C between atoms of the (0,1) and (1,1) shells; in D and E between atoms of the (1,1) and (1,2) shells.
Figure S7: Growth simulations of alkali metal clusters.
A Deposition of Cs atoms on a Na@Rb seed of 249 atoms, in which a Mackay Na inner core is covered by two Ch1 shells of Rb atoms. The growth temperature is 125 K and the deposition rate is 1 atoms/ns. Cs atoms for a Ch2 shell. B Deposition of K atoms on a Na@K seed of 45 atoms, in which a Mackay core of 13 atoms is covered by one Ch1 shell of K atoms. The growth temperature is 125 K and the deposition rate is 1 atoms/ns. Deposited K atoms form two further Ch1 shells. C Deposition of Rb atoms on a Na@K seed of 117 atoms, made of a Mackay inner core of 13 Na atoms and two Ch1 shells of K atoms. Growth temperature is 125 K and deposition rate is 1 atoms/ns. Deposited Rb atoms first form the (2,2) BG shell, which is completed at 239 atoms, and then the growth continues by the formation of a (2,3) Ch2 shell. Upon deposition of further Rb atoms, the (2,2) shell suddenly transforms into a Ch1 (1,3) shell, still covered by a (2,3) Ch2 shell, which then transforms into a Ch1 (1,4) shell to finally reproduce the natural path which better corresponds to the optimal size mismatch between atoms of the different shells.
Figure S8: Growth sequences for Ag atoms on Ni and on Co seeds.
Figure S9: Growth sequence for Ag atoms on a Cu seed.
Figure S10: Growth sequence for AgPdNi clusters.
A Deposition of Pd atoms on a Ni147 Mackay seed. Temperature 400 K, deposition rate 0.1 atoms/ns. An AM shell is formed. B Ag on a Ni147@Pd132 seed terminated by an AM1 shell. Temparature 300 K, deposition rate 0.1 atoms/ns. Ag atoms form an AM2 shell and then a defective BG shell. On top of the BG shell, a defective Ch3 shell grows.
Figure S11: Growth of Ag atoms on a Ni defective seeds.
The growth simulations are at $T=450$ K, and deposition rate 0.1 atoms/ns. At variance with Figs. SS8 and SS9, here the images represent different shells of the final structures, obtained at the end of the growth, instead of snapshots taken during the growth. In A growth starts from a Ni289 seed, which is terminated by a slightly incomplete (0,4) MC shell. in th efinal structure, the (0,4) shell is completed by Ag atoms, then covered by Ch1 shells. In B growth starts from a Ni329 seed, which is terminated by a small fragment of a (0,5) MC shell. in the final structure, that shell is completed by Ag atoms, which transform it into a Ch1 (1,4) shell, which is covered by further Ch1 shells. In C the growth starts from a Ni227 seed, which is terminated by a half (0,4) MC shell. in the final structure, that shell is completed by Ag atoms, so that a the resulting shell presents a Ni and an Ag half. On the Ni side, the shell is MC (0,4), while on the Ag side the shell is Ch1 (1,3). This is indicated by the bifurcation in the path. Further shells are of class Ch1.
Figure S12: Growth mechanism on AM shells.
A The four orange atoms identify a fourfold adsorption site on the (1,3) AM shell. B A deposited adatom (cyan) is preferentially trapped in a site of this type. C A new fourfold adsorption site is created on top for the four orange atom so that D it can trap a further deposited atom. E-G The process is self-replicating, finally leading to the formation of the (2,3) AM shell. H The atoms of (1,4) AM shell on top of a (1,3) AM shell find only threefold adsorption sites, identified by the orange atoms. Sites at the facet edge are twofold (see the blue atoms) so that they are not even locally stable for isolated adatoms.
@@
In the initial configurations, these clusters always contain a core with MC shells (0,0) and (0,1) made of Na atoms, while the other shells are made of K atoms. For MC@BG, the (1,1) BG shell is added to the core, whereas for MC@BG-Ch1 a further (1,2) shell is added (see Fig. S6)a). For MC@MC, the (0,2) shell is added to the core, whereas for MC@MC-MC a further (0,3) shell is added. Exchanges of type 1 and 2 (see Fig. S6b,c) are made on the initial configuration. The energy differences (in eV) between the configurations after and before the exchange of the atomic pair are reported. The exchange is favourable for negative differences, unfavourable otherwise.
@@
In the initial configurations, these clusters always contain a core with MC shells (0,0) and (0,1) made of Na atoms, while the third and fourth shells are made of K and Rb atoms, respectively. For MC@BG@Ch1, the (1,1) and the (1,2) shells are added to the core (see Fig. S6)a), whereas for MC@MC@MC the (0,2) and (0,3) shells are added. Exchanges of types 1, 2, 3 and 4 (see Fig. S6b-e) are made on the initial configuration. The energy differences (in eV) between the configurations after and before the exchange of the atomic pair are reported. The exchange is favourable for negative differences, unfavourable otherwise.

### v2

@@ Acknowledgments
Figure 9: Facet of an anti-Mackay shell.
Generalised anti-Mackay (AM) shells are non-closed-packed, achiral icosahedral shells. Each AM shell is identified univocally by the two AM indexes $p$ and $q$: $p$ is the number of particles on the side of the closed-packed equilateral triangle concentric to the icosahedral facet; $q$ is the number of particles on each median of the icosahedral facet, between the vertex of the icosahedral facet and the vertex of the inner triangle (see Figure 6a in the main text). As for the CK shells, simple geometric considerations allow us to identify the rules for packing AM shells into icosahedra, and to associate the optimal size mismatch to each possible interface between nearby shells of different kind.
Here we start by calculating the number of particle in a AM shell, as a function of the two AM indexes $p$ and $q$. It is useful to divide the particles of the icosahedral facet into different groups, as shown in Supplementary Figure 9:
@@
so that each AM shell can be put in relation with a CK shell of the same size. More precisely, the AM shell can transform into the corresponding CK shell by some concerted rotations of the particles, with different rotation axes (perpendicular to the planes of the icosahedral facets) and different rotation angles and, viceversa, CK shells can transform into AM ones. In Figure 6b-c of the main text we have shown two examples:
@@
The edge of the AM shell with unit inter-particle distance and AM indexes $p$, $q$ can be easily calculated by looking at Supplementary Figure 9. It corresponds to the side of the inner triangle, which is $p-1$, plus twice the projection of the segment between the vertices of the facet and of the inner triangle (which is $q+1$ long) on the facet edge:
@@
Values are reported in Figure 7a of the main text.
@@
The data on the optimal mismatch in Lennard-Jones and Morse clusters have been obtained by calculating the energy, after local minimization, as a function of the mismatch for the different types of structures, and looking for the mismatch at which the energy is minimum. Two examples are reported in Supplementary Figure 10, where we consider MC, AM1 and Ch1 shells on Mackay cores made of 5 and 8 MC shells. The MC structures are better stabilised in a low-mismatch range. The AM1 shells find their optimal mismatch in an intermediate range and after a threshold mismatch they become unstable, i.e. they are not even local minima. On the contrary, the Ch1 structures are stable in a high-mismatch range and become unstable below a threshold mismatch. In the cases reported in Supplementary Figure 10 the range of mismatch values in which both AM1 and Ch1 structures are local minima is very narrow. In Supplementary Figure 10a we have reported the same data of Figure 3a of the main text for the addition of either a MC or a Ch1 shell on a five-shell Mackay core; data for the addition of a AM1 shell have been included, in order to compare the three possible arrangements of the outer shell.
Figure 10: Stability of MC, AM1 and Ch1 shells in Lennard-Jones clusters.
Figure 11: Optimal mismatch for two AM shells on a MC core.
Figure 12: Stability of two AM shells on a MC core.
@@
Values of the optimal mismatch are reported in Supplementary Figure 11. As in the case of chiral shells (see Figure 3d in the main text), the optimal mismatch is larger for the class-changing step $(1,k)^{*}\to(2,k)^{*}$. This may lead to the conclusion that the outer shell should be preferentially of AM1 type (i.e. $(1,k+1)^{*}$) if the mismatch is zero or low. But this is not the case, as we show in Supplementary Figure 12. In fact the mismatch range in which the configuration with two AM1 shells is a local minimum is very narrow, and even when locally stable, its energy is quite high. Outside that range, the outer shell spontaneously reconstructs upon local minimization, as shown in Supplementary Figure 12a-b. This is due to the very unfavourable coordination of particles on the facet edges of the outer shell, which sit on only two nearest neighbours of the shell below (this point will be further discussed in the last Section). Both reconstructions improve that coordination by displacing atoms on fourfold sites of the substrate.
In Supplementary Figure 12c we calculate the energies of unreconstructed and reconstructed configurations of Lennard-Jones clusters. The reconstruction of Supplementary Figure 12b is more favourable, as it displaces a larger number of particles on sites of higher coordination with the substrate. Due to the instability of the configuration with two AM1 shells, the natural growth sequences lead to the formation of an outer AM2 shell, as shown in Figure 7b-c of the main text and discussed in the Supplementary Note 3.
Finally, we note that the configuration with two Ch1 shells, which has the same number of particles, is much more favourable in a wide range of mismatch, as shown in Supplementary Figure 12c. In fact, from the results in Supplementary Figure 12c it turns out also that the stability range of two Ch1 shells is enlarged with respect to that of one Ch1 shell (see Supplementary Figure 10a).
@@
Figure 13: Clusters of the DFT calculations.
@@
The structures considered in the calculations are of two types, with either $i=3$ or $i=4$ shells. In this figure we show only those containing BG and Ch1 shells, while the analogous structures containing 3 or 4 MC shells are not shown. The first structure ($i=3$) contains the central atom (0,0), the (0,1) MC shell, and the (1,1) BG shell, corresponding to the left and middle panels in a, for a total of 45 atoms. The second structure ($i=4$) contains in addition the (1,2) Ch1 shell, as shown in the right panel of a, for a total of 117 atoms. In b-e the different types of atom pair swaps are shown: in b and c between atoms of the (0,1) and (1,1) shells; in d and e between atoms of the (1,1) and (1,2) shells.
In Supplementary Tables 2-4 we report the data for the atomic pair exchanges shown in Supplementary Figure 13 for different types of alkali metal clusters. DFT [51] and Gupta potential [52] results are reported. The results have been obtained as explained in the Methods section of the main text. Here we recall that the energy differences between the configuration after and before the exchange are related to locally minimized configurations.
All structures contain a core made of $(0,0)$ and $(0,1)$ MC shells, upon which further shells are added, either of Ch1 class (the $(1,1)$ BG shell and the $(1,2)$ chiral shell) or of Ch0 class (the $(0,2)$ and $(0,3)$ MC shells). In the latter case, standard Mackay icosahedra are obtained. According to the estimates of the optimal mismatch values, for Na@K, Na@Rb and Na@K@Rb we expect that the clusters with Ch1 shells are more stable than the Mackay clusters. This is confirmed by the data reported in Supplementary Tables 2-4, where it is shown that Mackay clusters are energetically unstable upon atomic pair exchanges, while the clusters with Ch1 shells are stable. We note that DFT and Gupta results are in good agreement, which supports the validity of the Gupta model for these clusters.
@@
In the initial configurations, these clusters always contain a core with MC shells $(0,0)$ and $(0,1)$ made of Na atoms, while the other shells are made of K atoms. For MC@BG, the $(1,1)$ BG shell is added to the core, whereas for MC@BG-Ch1 a further $(1,2)$ shell is added (see Supplementary Figure 13a). For MC@MC, the $(0,2)$ shell is added to the core, whereas for MC@MC-MC a further $(0,3)$ shell is added. Exchanges of type 1 and 2 (see Supplementary Figure 13b,c) are made on the initial configuration. The energy differences (in eV) between the configurations after and before the exchange of the atomic pair are reported. The exchange is favourable for negative differences, unfavourable otherwise.
@@
In the initial configurations, these clusters always contain a core with MC shells $(0,0)$ and $(0,1)$ made of Na atoms, while the third and fourth shells are made of K and Rb atoms, respectively. For MC@BG@Ch1, the $(1,1)$ and the $(1,2)$ shells are added to the core (see Supplementary Figure 13a), whereas for MC@MC@MC the $(0,2)$ and $(0,3)$ shells are added. Exchanges of types 1, 2, 3 and 4 (see Supplementary Figure 13b,e) are made on the initial configuration. The energy differences (in eV) between the configurations after and before the exchange of the atomic pair are reported. The exchange is favourable for negative differences, unfavourable otherwise.
Appendix S3 Growth simulations
Here we present additional results on the simulations of the growth after deposition of atoms one-by-one on a preformed seed [44, 24]. These simulations show a few more examples of growth sequences that can be represented by paths in the hexagonal lattice.
In Supplementary Figure 14, three growth sequences for binary and ternary alkali metal clusters are shown. In Supplementary Figure 14a, we report the deposition of K atoms on a Na@K cluster terminated by a BG $(1,1)$ shell of K atoms. This is a case of deposition without mismatch, therefore the growth continues within the Ch1 class as expected.
In Supplementary Figure 14b, we show the result of the deposition of Cs atoms on a Na@Rb seed. The seed is terminated by a Ch1 $(1,3)$ shell of Rb atoms. According to the path rules, either a $(1,4)$ or a $(2,3)$ shell can form upon deposition of further atoms. The optimal size mismatch of these steps can be estimated by Eq. S53, where the coefficient $\xi$ is calculated recursively by applying Eq. S67 twice, since there are three shells made of the same atom type ($(1,1)$, $(1,2)$ and $(1,3)$). The optimal values of the mismatch are 0.0547 and 0.1162 for the steps towards $(1,4)$ and $(2,3)$, respectively. As usual, the class-conserving step is associated with the lower optimal mismatch. The mismatch between Cs and Rb, as evaluated from their bulk lattice constant, is of 0.077. Since Cs atoms are too large for accommodating on the $(1,4)$ shell, we expect the $(2,3)$ shell to form. This prediction finds confirmation in our growth simulations, where we observe the formation of a perfect Ch2 shell. We note that Cs and Rb atoms show some tendency towards intermixing [53], so that there are a few exchanges of Rb and Cs atoms during growth that bring some Rb atoms to the cluster surface.
In Supplementary Figure 14c a more complex growth sequence is shown. The seed is Na@K terminated by a Ch1 $(1,2)$ shell of K atoms, on which Rb atoms are deposited. Two steps are possible, leading to the formation of either the Ch1 $(1,3)$ shell or the BG $(2,2)$ shell. The optimal values of the mismatch are 0.0618 and 0.1129, respectively. The mismatch between K and Rb is 0.0716, which is slightly larger than the optimal mismatch for the growth of a Ch1 shell. Indeed, as in the case of Supplementary Figure 14b, the chirality class is changed and the BG $(2,2)$ shell is formed. This is followed then by a $(2,3)$ Ch2 shell, i.e. the only viable alternative for the growth on top of the BG shell. However this is not the end of the story, because upon further Rb atom deposition there is first a sudden transformation of the subsurface shell from $(2,2)$ to $(1,3)$ and then of the $(2,3)$ shell from $(2,3)$ to $(1,4)$. These transformations take place by sudden collective reshaping processes, of the type frequently observed in metal clusters [24]. The final result is as if the cluster would have grown into the Ch1 class from the beginning. The transformation of the first Rb shell from $(2,2)$ to $(1,3)$ is due to the packing of multiple shells, which affects the optimal mismatch at the Na-Rb interface. Specifically, the addition of further shells slightly increases the optimal mismatch (see Supplementary Note 1.7). The optimal mismatch for $(1,2)$-$(1,3)$ interface is therefore likely to approach and eventually overcome the mismatch between K and Rb, so that the Ch1 shell becomes more favourable than the BG one. This induces the reshaping of the shell. After the transformation of the first Rb shell, we have a $(2,3)$ Rb shell on top of a $(1,3)$ shell made of atoms of the same type. This configuration is not the optimal one, because when the mismatch is zero atoms are better accommodated on shells of the same chirality class. Therefore, the transformation of the second Rb shell towards the Ch1 $(1,4)$ arrangement takes place.
In Supplementary Figure 15 we show a growth sequence for a alkali metal cluster made of four different elements. The seed is Na@K@Rb terminated by a Ch1 $(1,3)$ shell of Rb atoms, on which Cs atoms are deposited. Two steps are possible, leading to the formation of either the Ch1 $(1,4)$ shell or the Ch2 $(2,3)$ shell. The optimal values of the mismatch are 0.0183 and 0.0778, respectively. We note that these values are different from those of Supplementary Figure 14, referring to the same steps on the closed-packed plane, because here in the seed there is a single shell of Rb atoms instead of three. The mismatch between Rb and Cs is 0.0769, which is very close to the optimal mismatch for the formation of the Ch2 $(2,3)$ shell. Such shell is indeed formed in our simulations. However, the deposition of further Cs atoms leads to the formation of the BG $(3,3)$ shell instead of the Ch2 $(2,4)$ shell, which is what one would expect when atoms of the same species are deposited. However, we note that, in this case, the optimal mismatch for both steps is quite small, being 0.0249 for the class-conserving step and 0.0459 for the class-changing one. As such, even though the Ch2 shell is expected to be more favourable, the BG one is not that unfavourable and therefore may form during the growth simulation. Besides, due to the mixing of Cs atoms with the smaller Rb atoms in the $(2,3)$ shell, the effective mismatch is actually larger then zero, which helps in stabilizing the BG shell. Upon further deposition of Cs atoms, the Ch3 $(3,4)$ shell is formed, corresponding to the natural growth step on top of the BG $(3,3)$ shell.
In Supplementary Figure 16 data on the growth of Ag atoms on Ni and Co Mackay seeds of 147 and 309 atoms are shown. The large mismatch (0.161 for Ag on Ni and 0.156 for Ag on Co) is close to optimal for the growth of Ch1 shells on seeds in this size range; this kind of growth is indeed observed in the growth simulations.
In Supplementary Figure 17 the growth of Ag atoms on a Mackay seeds of size 309 is shown. This cluster is terminated by a $(0,4)$ shell. For the growth of the first Ag shell, the mismatch between Ag and Cu (0.129) is closer to the optimal one of an AM1 shell. The AM1 $(1,4)^{*}$ shell is indeed grown, with very few defects. Then the growth on the AM1 shell continues by a class-changing path by forming defective AM shells up to the AM3 $(1,4)^{*}$ shell. Upon further deposition, there is a sudden transformation of all Ag shells to the Ch1 arrangement, up to the $(1,6)$; this configuration is energetically more favourable.
In Supplementary Figure 18 the growth of Ag on imperfect Mackay Ni seeds is shown. These seeds present an incomplete Mackay external shell. In spite of the defective character of the seed, Ch1 shells are formed. These shell are slightly distorted because of the asymmetries in the starting core, but they very much resemble the perfect Ch1 shells. This shows that the growth mode of chiral shell is quite robust against imperfections.
In Supplementary Figure 19 the growth mechanism of AM shells one on top of the other is discussed. This mechanism is dominated by the trapping of atoms in fourfold adsorption sites on the surface, which naturally leads to class changes at every step in the hexagonal plane.
Figure 14: Growth simulations of alkali metal clusters of two and three different elements.
a Deposition of K atoms on a Na@K seed of 45 atoms, in which a Mackay core of 13 atoms is covered by one Ch1 shell of K atoms. The growth temperature is 125 K and the deposition rate is 1 atoms/ns. Deposited K atoms form two further Ch1 shells. b Deposition of Cs atoms on a Na@Rb seed of 249 atoms, in which a Mackay Na inner core is covered by two Ch1 shells of Rb atoms. The growth temperature is 125 K and the deposition rate is 1 atoms/ns. Cs atoms form a Ch2 shell. c Deposition of Rb atoms on a Na@K seed of 117 atoms, made of a Mackay inner core of 13 Na atoms and two Ch1 shells of K atoms. Growth temperature is 125 K and deposition rate is 1 atoms/ns. Deposited Rb atoms first form the $(2,2)$ BG shell, which is completed at 239 atoms, and then the growth continues by the formation of a $(2,3)$ Ch2 shell. Upon deposition of further Rb atoms, the $(2,2)$ shell suddenly transforms into the $(1,3)$ Ch1 shell, still covered by a $(2,3)$ Ch2 shell, which then transforms into the $(1,4)$ Ch1 shell.
Figure 15: Growth simulations of alkali metal clusters of four different elements.
Deposition of Cs atoms on a Na@K@Rb seed of 249 atoms, in which a Mackay core of 13 atoms is covered by two Ch1 shell of K atoms, and by a Ch1 shell of Rb. The growth temperature is 125 K and the deposition rate is 1 atoms/ns. Deposited Cs atoms form Ch2 and Ch3 shells
Figure 16: Growth sequences for Ag atoms on Ni and on Co seeds.
Figure 17: Growth sequence for Ag atoms on a Cu seed.
Figure 18: Growth of Ag atoms on a Ni defective seeds.
The growth simulations are at $T=450$ K, and deposition rate 0.1 atoms/ns. At variance with Supplementary Figures 16 and 17, here the images represent different shells of the final structures, obtained at the end of the growth, instead of snapshots taken during the growth. In a growth starts from a Ni289 seed, which is terminated by a slightly incomplete $(0,4)$ MC shell. In the final structure, the $(0,4)$ shell is completed by Ag atoms, then covered by Ch1 shells. In b growth starts from a Ni329 seed, which is terminated by a small fragment of a (0,5) MC shell. in the final structure, that shell is completed by Ag atoms, which transform it into a Ch1 $(1,4)$ shell, which is covered by further Ch1 shells. In c the growth starts from a Ni227 seed, which is terminated by a half $(0,4)$ MC shell. In the final structure, that shell is completed by Ag atoms, so that a the resulting shell presents a Ni and an Ag half. On the Ni side, the shell is MC $(0,4)$, while on the Ag side the shell is Ch1 $(1,3)$. This is indicated by the bifurcation in the path. Further shells are of class Ch1.
Figure 19: Growth mechanism on AM shells.
a The four orange atoms identify a fourfold adsorption site on the $(1,3)^{*}$ AM shell. b A deposited adatom (cyan) is preferentially trapped in a site of this type. c A new fourfold adsorption site is created on top for the four orange atom so that d it can trap a further deposited atom. e-g The process is self-replicating, finally leading to the formation of the $(2,3)^{*}$ AM shell. h The atoms of $(1,4)^{*}$ AM shell on top of a $(1,3)^{*}$ AM shell find only threefold adsorption sites, identified by the orange atoms. Sites at the facet edge are twofold (see the blue atoms) so that they are not even locally stable for isolated adatoms.